OpenStudy (anonymous):

Simplify when it is possible. 10x^2 -y + x^2 =

OpenStudy (anonymous):

\[10x^2 -y + x^2 = \]

OpenStudy (fibonaccichick666):

DO you have any like terms?

OpenStudy (anonymous):

\[x^2\]

OpenStudy (fibonaccichick666):

yup so just combine those

OpenStudy (anonymous):

take all x terms on one side.. and all x terms on another.. u r done.. and 10+1=11

OpenStudy (anonymous):

x^4

OpenStudy (anonymous):

so its 11??

OpenStudy (fibonaccichick666):

no, you don't have to multiply notice, that the terms are just like your 4ab and 3ab from last time

OpenStudy (anonymous):

ok x^2 will cancel your left with 10+y right so its 10y....???

OpenStudy (fibonaccichick666):

only this time we have \(10x^2\) and \((1)x^2\)

OpenStudy (fibonaccichick666):

no, ignore what that other guy said

OpenStudy (anonymous):

so its 11Y

OpenStudy (anonymous):

im so confused

OpenStudy (fibonaccichick666):

no, you cannot combine x and y. They are different variables

OpenStudy (anonymous):

okay im asking what is the final answer

OpenStudy (fibonaccichick666):

I'm making you figure it out. what is 3b+4b=?

OpenStudy (anonymous):

7b^2

OpenStudy (fibonaccichick666):

no, you don't multiply the variable. just think if i had 3b+4b that is the same as (b+b+b)+(b+b+b+b)=?

OpenStudy (anonymous):

so then it just 7b

OpenStudy (anonymous):

omg im so confused

OpenStudy (fibonaccichick666):

yup, now for ours we have 10x^2 that is the same as \([x^2+x^2+x^2+x^2+x^2+x^2+x^2+x^2+x^2+x^2]+[x^2]=?\)

OpenStudy (anonymous):

ty i got it now

OpenStudy (fibonaccichick666):

ok, thanks for working through it